OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 8x - (5x + 4) greater than or equal to -31? x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left. A number line with a closed circle on -9 and shading to the right.

OpenStudy (anonymous):

what are the "following"?

OpenStudy (anonymous):

x greater than or equal to -9 -9 less than or greater to x A number line with a closed circle on -9 and shading to the left. A number line with a closed circle on -9 and shading to the right.

OpenStudy (anonymous):

I have no idea where to start

OpenStudy (texaschic101):

8x - (5x + 4) >= -31 8x - 5x - 4 >= -31 3x - 4 >= -31 3x >= -31 + 4 3x > = - 27 x >= - 9 Quick lesson : If there is an equal sign involved, it will have a closed circle If no equal sign is involved,the circle will be open. Greater then mans shading to the right Less then means shading to the left So what do you think the answer is ?

OpenStudy (texaschic101):

oops..typo...* means

OpenStudy (anonymous):

So the first one, x is greater than or equal to -9?

OpenStudy (texaschic101):

you are looking for the one that does not represent the solution. The first answer does represent the solution. Try again.

OpenStudy (anonymous):

Oops,my bad. So it would be the last one.

OpenStudy (texaschic101):

pay attention to your inequality sign....if it is greater, shading goes to the right, if it is less then shading goes to the left. Is the sign greater or less then ?

OpenStudy (anonymous):

I always thought whichever way the sign pointed was the way you shaded. So < shades left, and vice versa. The sign is greater than, x is greater than or equal to -9

OpenStudy (texaschic101):

< shades to the left and > shades to the right. The inequality sign in this problem is greater,therefore shading to the right is a correct solution. However, shading to the left is not .

OpenStudy (anonymous):

I keep forgetting it says which one is NOT. Algebra is definitely a weak spot I have, haha

OpenStudy (anonymous):

Thank you so much for your help!

OpenStudy (texaschic101):

I hate geometry....algebra is ok. Glad I could help :)

OpenStudy (anonymous):

I have to take geometry this coming school year D: